OpenStudy (anonymous):

perform the indicated operation 8/9-(-1/2) help!!!

jimthompson5910 (jim_thompson5910):

8/9-(-1/2) is the same as 8/9+1/2

jimthompson5910 (jim_thompson5910):

From there, you need to find the LCD and get every denominator equal to the LCD

jimthompson5910 (jim_thompson5910):

once you've done that, you can add

OpenStudy (anonymous):

ok this what i got but doesnt look right 8/9*18/18+1/2*18/18=144/162+18/36=162/198? what am i doing wrong?

jimthompson5910 (jim_thompson5910):

close, but you don't multiply by the LCD

OpenStudy (anonymous):

ok i add?

jimthompson5910 (jim_thompson5910):

8/9 multiply top and bottom by 2/2 to get (8*2)/(9*2) 16/18 So 8/9 = 16/18

jimthompson5910 (jim_thompson5910):

then multiply 1/2 by 9/9 to get 9/18 So 8/9+1/2 becomes 16/18 + 9/18

OpenStudy (anonymous):

oh forgot to flip it to 2*one side an 18*the other!!!thx

jimthompson5910 (jim_thompson5910):

yw